mommymeanna16:

I need help but there are a lot of problems if sum 1 is able to stay and help that would be great ss below

mommymeanna16:

Buenos88:

u=2 6*6=36

mommymeanna16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Buenos88 u=2 6*6=36 \(\color{#0cbb34}{\text{End of Quote}}\) thank u do u think u could help me with more if not thats okay

Buenos88:

Yeah, why don't you just dm me?

mommymeanna16:

ok

itsmehjay:

\(\color{#0cbb}{\text{Originally Posted by}}\) @Buenos88 u=2 6*6=36 \(\color{#0cbb}{\text{End of Quote}}\) You can't give direct answers

itsmehjay:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @itsmehjay \(\color{#0cbb}{\text{Originally Posted by}}\) @Buenos88 u=2 6*6=36 \(\color{#0cbb}{\text{End of Quote}}\) You can't give direct answers \(\color{#0cbb34}{\text{End of Quote}}\) ok.. now I'm not sure. well does this count as a direct answer? @carmelle @IhelpUuHELPme

IhelpUuHELPme:

yes, that is considered a direct answer bc you didn't explain why u equals 2, and such as the steps too.

itsmehjay:

Yeah what she said >:T So no more of that bub

Buenos88:

ok